(a) Solve for 0 < θ < 5, the equation 5+3cosec (20θ+1) = 0


(b) FInd all the angles between -1800 and 1800 such that


                  cosec y + 11cot y = 4sin y


Thanks,


Vishnu

Grade:10

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Vishnu

a )

5+3cosec (20θ+1) = 0

sin(20θ+1)=-3/5

0 < θ < 5

so  1 < 20 θ +1< 101

for this angel range sin (20θ+1) will always be positive so no solution

b)

cosec y + 11cot y = 4sin y

simplify    4cos2y +11cosy -3=0

 this give cos y=-7/4 and 3/2   which is not posible so no solution exits


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin